# Multi-Sampling-Frequency Naturalness MOS Prediction Using Self-Supervised Learning Model with Sampling-Frequency-Independent Layer

### v1

@@ Abstract
In this paper, we present “MSR-UTMOS”, our MOS prediction model capable of handling speech with multiple SFs, developed for the AudioMOS Challenge (AMC) 2025 Track 3. This track aimed to accurately predict the results of naturalness MOS test conducted on synthetic speech at 16, 24, and 48 kHz SFs. Our core contribution is the integration of an SF-independent (SFI) convolutional layer [12] into an SSL model. We call this architecture “SFI-SSL model.” The SFI layer adjusts its weights in accordance with the SF of the input signal, enabling a single MOS prediction model to process speech at various SFs, including those utilizing frequency components higher than the trained Nyquist frequency. Our submission to the AMC 2025 Track 3 ranked the first in utterance-level mean squared error (MSE) and the fourth in the final ranking based on system-level Spearman’s rank correlation coefficient (SRCC). We also report results of an ablation study to investigate essential factors in our model: SFI-SSL, knowledge distillation (KD) for initializing the parameters of an SFI-SSL model, and fine-tuning the model on MOS prediction. The code and pretrained SFI-SSL models are available online.
@@ III-B SFI Convolutional Layer
The SFI convolutional layer is an extension of the convolutional layer for handing arbitrary SFs. This layer is based on the fact that the weights of the convolutional layer can be interpreted as the collection of digital filters. From this interpretation, this layer generates the weights from a function representing the frequency response, which we call the latent analog filter. By utilizing digital filter design method, the generated weights are consistent for any SFs. Thus, this layer can handle arbitrary SFs. See [12] for details of the digital filter design method. Fig. 2 shows the schematic of the SFI convolutional layer. For the latent analog filter, we use neural analog filter (NAF), which represents the frequency response via a DNN [13]. The NAF consists of an input transformation method based on random Fourier features (RFFs) [14] and a simple feedforward network.
@@ IV-B Experimental conditions
Datasets: For KD-based initialization of the proposed SFI-SSL models, we used all speech samples from EARS [16], a 100 hours of 48 kHz sampled English speech by 107 speakers. For fine-tuning the non-SFI- and SFI-SSL models on the MOS prediction, we used the train subset of BVCC [17] containing the naturalness scores of 4,974 synthesized speech. Each of the samples is associated with naturalness ratings from 8 listeners, thus the total number of ratings is 39,794 ratings. For the final training, we used the AMC 2025 Track 3 official training/validation sets.
@@
DNN optimization: We used schedule-free version of RAdam optimizer [22, 23] with its hyperparameters setting to , , and , which is available on GitHub. We set the batch size to 32 for the KD, 64 for the pretraining, and 24 for the final training. The KD was performed for one day with single H200 GPU. The pretraining and final training were performed for 50 epochs, which was equivalent to 8 hours with 1/7-th of H200 GPU. We selected the best checkpoints based on the MSE on the validation set of BVCC for pretraining. For the final training, the checkpoint with minimum loss on the validation set acquired by cross validation were selected. Note that this validation set was different from the Track 3 official validation set.
@@ V Conclusion
Acknowledgements: This work was supported by JST Moonshot JPMJMS2011, JST BOOST JPMJBY24C9, and JSPS KAKENHI JP23K28108.

### v2

@@ Abstract
In this paper, we present “MSR-UTMOS”, our MOS prediction model capable of handling speech with multiple SFs, developed for the AudioMOS Challenge (AMC) 2025 Track 3. This track aimed to accurately predict the results of naturalness MOS test conducted on synthetic speech at 16, 24, and 48 kHz SFs. Our core contribution is the integration of an SF-independent (SFI) convolutional layer [12] into an SSL model. We call this architecture “SFI-SSL model.” The SFI layer adjusts its weights in accordance with the SF of the input signal, enabling a single MOS prediction model to process speech at various SFs, including those utilizing frequency components higher than the trained Nyquist frequency. Our submission to the AMC 2025 Track 3 ranked the first in utterance-level mean squared error (MSE) and the fourth in the final ranking based on system-level Spearman’s rank correlation coefficient (SRCC). We also report results of an ablation study to investigate essential factors in our model: SFI-SSL, knowledge distillation (KD) for initializing the parameters of an SFI-SSL model, and fine-tuning the model on MOS prediction. The code111https://github.com/sarulab-speech/sfi-utmos and pretrained SFI-SSL models222https://huggingface.co/collections/Wataru/sfi-ssl-6852d2a9382ab3c8b65fc57c are available online.
@@ III-B SFI Convolutional Layer
The SFI convolutional layer is an extension of the convolutional layer for handling arbitrary SFs. This layer is based on the fact that the weights of the convolutional layer can be interpreted as the collection of digital filters. From this interpretation, this layer generates the weights from a function representing the frequency response, which we call the latent analog filter. By utilizing digital filter design method, the generated weights are consistent for any SFs. Thus, this layer can handle arbitrary SFs. See [12] for details of the digital filter design method. Fig. 2 shows the schematic of the SFI convolutional layer. For the latent analog filter, we use neural analog filter (NAF), which represents the frequency response via a DNN [13]. The NAF consists of an input transformation method based on random Fourier features (RFFs) [14] and a simple feedforward network.
@@ IV-B Experimental conditions
Datasets: For KD-based initialization of the proposed SFI-SSL models, we used all speech samples from EARS [16], a 100 hours of 48 kHz sampled English speech by 107 speakers. For fine-tuning the non-SFI- and SFI-SSL models on the MOS prediction, we used the train subset of BVCC [17] containing the naturalness scores of 4,974 synthesized speech. Each of the samples is associated with naturalness ratings from 8 listeners, thus the total number of ratings is 39,794 ratings. For the final training, we used the AMC 2025 Track 3 official sets.
@@
DNN optimization: We used schedule-free version of RAdam optimizer [22, 23] with its hyperparameters setting to , , and . We set the batch size to 32 for the KD, 64 for the pretraining, and 24 for the final training. The KD was performed for one day with single H200 GPU. The pretraining and final training were performed for 50 epochs, which was equivalent to 8 hours with 1/7-th of H200 GPU. We selected the best checkpoints based on the MSE on the validation set of BVCC for pretraining. For the final training, the checkpoint with minimum loss on the validation set acquired by cross validation were selected. Note that this validation set was different from the Track 3 official validation set.
@@ V Conclusion
Acknowledgements: This work was supported by JST Moonshot JPMJMS2011, JST BOOST JPMJBY24C9, JSPS KAKENHI JP23K28108 and Research Grant S of the Tateisi Science and Technology Foundation.